Find the slope of a line perpendicular to the line passing through the points
(-1,4) and (8,1)

Answer by checkley75(3666)

SLOPE=(Y2-Y1)/(X2-X1)
SLOPE=(1-4)/8+1)
SLOPE=-3/9
SLOPE=-1/3 SO A LINE PERPENDICULAR WILL HAVE A SLOPE OF 3.